OpenStudy (lilai3):

Please HELP!!! 22 GALLONS PER DAY EQUALS TO HOW MANY ML PER MINUTE???? LOLS sorry caps BUt plz help. plz

hartnn (hartnn):

ok, u know how 1 gallon = .... ml ?

hartnn (hartnn):

and how many minutes are there in a da ?

hartnn (hartnn):

*day

OpenStudy (lilai3):

1 gal= 785 mL

OpenStudy (lilai3):

1440 min=1 day

OpenStudy (lilai3):

mins*

hartnn (hartnn):

u sure about 1 gal= 785 mL ?

hartnn (hartnn):

1440 min=1 day is correct

OpenStudy (lilai3):

@hartnn, the website you gave me also says 1 gal= 3785 ML

hartnn (hartnn):

oh, u missed 3 before....

hartnn (hartnn):

so, 22 GALLONS/DAY = 22 * 3785 ML/DAY = 22*3785 / 1440 ML/MIN.

OpenStudy (lilai3):

oh oops didn't notice my mistake sorry

hartnn (hartnn):

no problem

OpenStudy (lilai3):

uhm can you like draw it out more clearly.. by draw i mean explain

OpenStudy (lilai3):

and the after part of the answer of what you get too please

hartnn (hartnn):

ok, 22 GALLONS = ?? ML

hartnn (hartnn):

if 1 GALLON= 3785 ML 22 GALLONS = ?? ML

OpenStudy (lilai3):

83270

hartnn (hartnn):

yes, so, 83270 ML/DAY = ?? ML/MIN

OpenStudy (lilai3):

uh... 1440?

hartnn (hartnn):

83270 ML/DAY = 83270 ML / (1440 MIN) u got this ?

hartnn (hartnn):

becuse, 1 day =1440 min

OpenStudy (lilai3):

yep i got that...wait i was right about the 1440? oh yay i just guessed LOLS

hartnn (hartnn):

ok, 83270 ML / (1440 MIN) = 83270/1440 ML/MIN ok ?

OpenStudy (lilai3):

ok.

hartnn (hartnn):

83270 ML / (1440 MIN) = 83270/1440 ML/MIN= ??? ML/MIN

hartnn (hartnn):

so what answer u got ?

OpenStudy (lilai3):

uhm... 58. sumthing

hartnn (hartnn):

yeah, 57.8 so its 57.8 ML PER MINUTE

OpenStudy (lilai3):

okay is that it that's the answer wow that seems so simple i'm a nincompoop. can you help me for 3 more problems if you do not mind?

hartnn (hartnn):

thats it. it was simple. new question, new post.